# Degree-one foliations on complete intersections

### v1

@@ Abstract.
We prove that, under mild restrictions, the space of codimension-one foliations of degree one on a smooth projective complete intersection has two irreducible components of logarithmic type. We also prove that the same conclusion holds for any smooth hypersurface of dimension at least three that is not a quadric threefold. The proof of these results follows essentially from a more general structure theorem for foliations on manifolds covered by lines.
@@ 1. Introduction
A singular holomorphic foliation $\mathcal{F}$ of codimension-one on a projective variety $X$ is determined by a $1$-form $\omega$ with values in a line bundle $N$, satisfying the Frobenius integrability condition $\omega\wedge d\omega=0$, and whose singular set has codimension at least two. In this case, $N$ is called the normal bundle of $\mathcal{F}$. The set of all such foliations with fixed normal bundle $N$ defines a quasi-projective variety, denoted by ${\sf Fol}(X,N)$, known as the space of codimension-one foliations on $X$ with normal bundle $N$.
The irreducible components of these spaces have been studied in various contexts. When $\operatorname{Pic}(X)\simeq\mathbb{Z}$, a foliation $\mathcal{F}$ with normal bundle $N$ is said to have degree $d$ if $N\simeq{\mathcal{O}}_{X}(d+2)$, where ${\mathcal{O}}_{X}(1)$ denotes the hyperplane bundle. In this case, we denote the space of degree-$d$ foliations by ${\sf Fol}(X,d)$. For example, when $X$ is a surface, the integrability condition is automatically satisfied, and the corresponding spaces are irreducible. When $X={\mathbb{P}^{n}}$, for $n\geq 3$, the cases $d=0,1,2$ are completely classified, and partial results exist for $d=3$ (see [13], [16], [7], and [10]). For more general smooth projective varieties, few results describe the structure of foliations of low degree. When $d=0$, classifications are available for certain homogeneous varieties [5, 19], for hypersurfaces [15], and for smooth complete intersections [1].
@@
Let $X$ be a smooth complete intersection in ${\mathbb{P}^{n+s}}$ of type $(d_{1},\ldots,d_{s})$, with $n\geq 4$ and $c_{1}(TX)\geq 2$. If either $s=1$ (i.e., $X$ is a hypersurface) or $d_{j}\geq 3$ for all $j=1,\dots,s$, then ${\sf Fol}(X,1)$ has exactly two irreducible components.
The general elements in these components are defined by logarithmic forms (see §2.2). In particular, all foliations in these spaces arise as restrictions of foliations on the ambient projective space (see Proposition 3.1). When $X$ is the smooth three-dimensional quadric, it is shown in [20] that ${\sf Fol}(X,1)$ possesses an additional irreducible component, consisting of rigid foliations arising from actions of the affine Lie algebra on the quadric, totaling three components.
Among hypersurfaces, the quadric is the only exceptional case. Theorem A applies to all smooth hypersurfaces of degree $\deg(X)\leq n$ and dimension $n\geq 4$. Moreover, by [15, Theorem 3.6], every degree-one foliation on a hypersurface $X$ extends to ${\mathbb{P}^{n+1}}$ when $n\geq 3$ and $\deg(X)\geq 4$. We complement this result by showing that it also holds for cubic threefolds:
@@
To prove Theorem A, we focus on varieties covered by lines. We study the deformations of lines along the leaves of a foliation, as introduced in [21]. First, we show that a general free line is not invariant under any foliation (see Proposition 4.7).
@@
Let $X\subset{\mathbb{P}^{N}}$ be a smooth projective variety of dimension $n\geq 3$ with $\operatorname{Pic}(X)=\mathbb{Z}$, covered by lines. If $\mathcal{F}$ is a codimension-one singular holomorphic foliation on $X$ of degree $d$ such that either $d=1$ or
@@
Specializing to degree-one foliations on smooth complete intersections with $c_{1}(TX)\geq 2$, we conclude that either the foliation is defined by a closed rational 1-form without codimension-one zeros, or it admits a codimension-two subfoliation of degree zero (see Corollary 4.9). In the latter case, we prove that the foliation is the pullback of a degree-one foliation on the projective plane, and is therefore again defined by a closed rational 1-form without codimension-one zeros. This completes the proof of Theorem A.
Finally, we apply the same techniques to analyze the structure of degree-two foliations on smooth quadrics (see Corollary 4.10 and Theorem 4.11), providing a partial classification for this setting and laying groundwork for future results.
@@ 2. Holomorphic Foliations
Let $X$ be a smooth projective complex manifold. A distribution $\mathcal{F}$ on $X$ is given by a saturated subsheaf ${T_{\mathcal{F}}}\subset TX$, often called the tangent sheaf of $\mathcal{F}$. Here, by saturated we mean that the quotient $TX/{T_{\mathcal{F}}}$ is a torsion-free sheaf. The singular locus of $\mathcal{F}$, denoted by $\operatorname{Sing}(\mathcal{F})$, is the locus where the quotient $TX/{T_{\mathcal{F}}}$ is not locally free. Since $TX/{T_{\mathcal{F}}}$ is torsion-free, it follows that $\operatorname{codim}\operatorname{Sing}(\mathcal{F})\geq 2$. We define the normal sheaf of $\mathcal{F}$ as ${N_{\mathcal{F}}}=(TX/{T_{\mathcal{F}}})^{\vee\vee}$.
@@
The map ${TX}\to{N_{\mathcal{F}}}$ is defined by contraction with a twisted $p$-form $\omega_{\mathcal{F}}\in H^{0}(X,\Omega^{p}_{X}\otimes\det({N_{\mathcal{F}}}))$, where $p=\operatorname{rk}{N_{\mathcal{F}}}$. Conversely, given a line bundle $L\in\operatorname{Pic}(X)$, $\omega\in H^{0}(X,\Omega^{p}_{X}\otimes L)$ defines a distribution if and only if it is locally decomposable off the singular set, LDS for short, see [11]. This condition means that for every point $x\in X\setminus\operatorname{Sing}(\omega)$ there exists an open neighborhood $U\ni x$ and $1$-forms $\alpha_{1},\dots,\alpha_{p}$ such that
@@
The main objective of our work is to describe the irreducible components of this space for $p=1$, $X\subset{\mathbb{P}^{n+1}}$ a smooth complete intersection and specified $L$. Any line bundle on ${\mathbb{P}^{n}}$ is isomorphic to ${\mathcal{O}}_{\mathbb{P}^{n}}(k)$ for some $k\in{\mathbb{Z}}$. A distribution has degree $d$ if it is defined by $\omega\in H^{0}(\Omega_{\mathbb{P}^{n}}^{p}(d+p+1))$. This follows Lins Neto’s convention of the degree of a foliation. In general, if $X$ is a smooth projective variety such that $\operatorname{Pic}(X)={\mathbb{Z}}$, we define the degree of a foliation $\mathcal{F}$ determined by a section in $H^{0}(X,\Omega_{X}^{p}(d+p+1))$ as $\deg(\mathcal{F}):=d$. In this case, for simplicity, we will denote ${\sf Fol}(X,d):={\sf Fol}^{1}(X,\mathcal{O}_{X}(d+2))$. We are particularly interested in smooth complete intersections.
@@
In [1, 14, 15] the authors study distributions and foliations on smooth complete intersections satisfying an extra hypothesis: $X=V(f_{1},\dots,f_{k})$ with $\deg(f_{1})\leq\dots\leq\deg(f_{k})$ such that $X_{i}=V(f_{1},\dots,f_{i})$ is smooth for every $i$. There are smooth complete intersections without this property; for instance, $X=V(f_{1},f_{2})$ where $\deg(f_{1})<\deg(f_{2})$ and $V(f_{1})$ is singular. However, one can verify that the arguments in these papers depend uniquely on the existence of a chain of smooth complete intersection $X_{1}\supset\cdots\supset X_{k}=X$, regardless of the degrees. Therefore, in light of Lemma 2.1, their results apply to smooth complete intersections in general.
@@
Note further that a smooth complete intersection only admits distributions and foliations of nonnegative degree. Indeed, $H^{0}(X,\Omega_{X}^{p}(d+p+1))=0$ for $d<0$ and $X$ a smooth complete intersection, see [1, Lemma 5.17].
@@ 2.1. Algebraically integrable foliations
A foliation on $X$ is algebraically integrable if its general leaf is an algebraic subvariety. In this case, there exists a dominant rational map $\pi\colon X\dashrightarrow Y$ onto a projective variety $Y$ such that the closure of a leaf is an irreducible component of a fiber of $\pi$, see [2, §3].
@@
Let $X$ be a complex projective manifold of dimension $n\geq 3$ with $\operatorname{Pic}(X)\cong{\mathbb{Z}}$. Let $\mathcal{F}$ be a codimension-one foliation on $X$. Suppose that $\mathcal{F}$ is algebraically integrable, then either:
@@
This follows directly from the arguments in [20, Lemma 3.1 and Theorem 3.2]. ∎
@@
Let $X\subset{\mathbb{P}^{N}}$ be a smooth complete intersection of dimension $n\geq 3$ and $\mathcal{F}$ be a codimension-one singular holomorphic foliation on $X$ of degree $d$. If $\mathcal{F}$ is algebraically integrable, then
@@
By Proposition 2.2, either $\mathcal{F}$ admits a rational first integral $\rho\colon X\dashrightarrow{\mathbb{P}^{1}}$ with only irreducible fibers or it admits a subfoliation $\mathcal{G}$. Suppose we are in the first case. By Lefschetz Hyperplane Theorem, $X$ is simply connected. Then the argument follows the proof of [20, Proposition 3.5]. By [20, Theorem 3.3], $\rho$ admits at most $2$ multiple fibers, which we may assume to be over 00 and $\infty$, and write $\rho^{-1}(0)=pH_{0}$ and $\rho^{-1}(\infty)=qH_{\infty}$. Then $\rho^{*}(\frac{dx}{x}-\frac{dy}{y})$ is a logarithmic $1$-form defining $\mathcal{F}$ with polar divisor $H_{0}+H_{\infty}$ and empty zero divisor. ∎
Notice that the proof of corollary holds for any simply connected manifold $X$ with Picard number $1$. For our purposes, we will only need $X$ to be a smooth complete intersection, especially hypersurfaces.
@@ 2.2. Irreducible components of the space of foliations
where $L_{1}$ and $L_{2}$ are degree one homogeneous polynomials, and thus they are all conjugate to each other (see [13, Proposition 3.1]). Therefore, ${\sf Fol}({\mathbb{P}^{n}},0)$ is isomorphic to the Grassmannian of lines in ${\mathbb{P}^{n}}$.
For codimension-one foliations of degree one, Jouanoulou in [16] proved that ${\sf Fol}({\mathbb{P}^{n}},1)$ has two irreducible components. Both components parameterize logarithmic foliations, which we now describe in more generality.
Let $X$ be any smooth projective variety such that $\dim(X)\geq 3$, $\operatorname{Pic}(X)={\mathbb{Z}}$ and $H^{1}(X,{\mathbb{C}})=0$. Given $d_{1},d_{2},\dots,d_{k}$ positive integers and $\lambda_{1},\lambda_{2},\dots,\lambda_{k}$ complex numbers, not all zero, such that $\sum\lambda_{i}\cdot d_{i}=0$, we can take sections $f_{i}\in H^{0}(X,{\mathcal{O}}_{X}(d_{i}))$ to form global sections
@@
Then $\omega$ is integrable and, if its singular locus has codimension greater than one, it determines a foliation known as a logarithmic foliation. Varying $\lambda_{i}$ and $f_{i}$, $i=1,\dots,k$, the forms $\omega$ sweep an irreducible algebraic subvariety of ${\sf Fol}(X,d_{1}+\dots+d_{k}-2)$ whose closure is denoted by $\textsf{Log}(X;d_{1},d_{2},\dots,d_{k})$. In [6], it was established that $\textsf{Log}(X;d_{1},d_{2},\dots,d_{k})$ is an irreducible component of ${\sf Fol}(X,\sum d_{i}-2)$. It is called the logarithmic component. If $k=2$, it is also called a rational component. These varieties can also be defined by the closure of the rational map
@@
where $\mathbb{L}_{2}$ is the set of linear projections of ${\mathbb{P}^{n}}$ to ${\mathbb{P}^{2}}$. Then, in [8], it was demonstrated that $\textsf{LPB}(d)$ is an irreducible component of ${\sf Fol}({\mathbb{P}^{n}},d)$, and it is called the Linear Pull-Back component. It is an elementary exercise to show that $\textsf{LPB}(1)=\textsf{Log}({\mathbb{P}^{n}};1,1,1)$.
For degree two, Cerveau and Lins Neto proved in [7] that ${\sf Fol}({\mathbb{P}^{n}},2)$ has six irreducible components: a pull-back Linear component $\textsf{LPB}(2)$; the rational components $\textsf{Log}({\mathbb{P}^{n}};3,1)$ and $\textsf{Log}({\mathbb{P}^{n}};2,2)$; the logarithmic components $\textsf{Log}({\mathbb{P}^{n}};2,1,1)$ and $\textsf{Log}({\mathbb{P}^{n}};1,1,1,1)$; and another irredutible component called exeptional component, which is determined by the closure of the map
@@
$\beta=\frac{1}{x_{0}}(2fdg-3gdf)\in H^{0}({\mathbb{P}^{3}},\Omega^{1}_{{\mathbb{P}^{3}}}(4)).$
where $f=2x_{0}x_{2}-x_{1}^{2}$ and $g=3x_{0}^{2}x_{3}-3x_{0}x_{1}x_{2}+x_{1}^{3}$. In [10], the irreducible components of ${\sf Fol}({\mathbb{P}^{n}},3)$ were partially classified; there are at least 24 irreducible components.
Besides the projective space, there is some literature on irreducible components of the space of foliations on, for instance, Fano three-folds [20], homogeneous spaces [5, 19], and hypersurfaces [1, 15].
In the case of smooth projective hypersurfaces $X$ in ${\mathbb{P}^{n+1}}$, $n\geq 3$, if
@@
then the restriction of foliations from the ambient ${\mathbb{P}^{n+1}}$ defines an isomorphism between ${\sf Fol}(X,d)$ and ${\sf Fol}({\mathbb{P}^{n+1}},d)$, see [15, Theorem 3.6]. Therefore, the previous classifications for foliations on projective spaces are valid for these varieties. The problem is then to study low-degree hypersurfaces; for degree-one foliations, we need to analyze hypersurfaces of degree at most three.
In [20], the authors proved that for $X\subset{\mathbb{P}^{4}}$ a smooth quadric hypersurface, ${\sf Fol}(X,1)$ has three irreducible components; only two are restrictions from foliations on ${\mathbb{P}^{4}}$. In this work, we prove that, for hypersurfaces, this is the only case where ${\sf Fol}(X,1)$ has more than two irreducible components.

### v2

@@ Abstract.
We prove that, under mild restrictions, the space of codimension-one foliations of degree one on a smooth projective complete intersection has two irreducible components of logarithmic type. We also prove that the same conclusion holds for any smooth hypersurface of dimension at least three that is not a quadric threefold. The proofs of these results follow essentially from a more general structure theorem for foliations on manifolds covered by lines.
@@ 1. Introduction
A singular holomorphic foliation $\mathcal{F}$ of codimension one on a projective variety $X$ is determined by a $1$-form $\omega$ with values in a line bundle $N$, satisfying the Frobenius integrability condition $\omega\wedge d\omega=0$, and whose singular set has codimension at least two. In this case, $N$ is called the normal bundle of $\mathcal{F}$. The set of all such foliations with fixed normal bundle $N$ defines a quasi-projective variety, denoted by ${\sf Fol}(X,N)$, known as the space of codimension-one foliations on $X$ with normal bundle $N$.
The irreducible components of these spaces have been studied in various contexts. For example, when $X$ is a surface, the integrability condition is automatically satisfied, and the corresponding spaces are irreducible. When $\operatorname{Pic}(X)\simeq\mathbb{Z}$, a foliation $\mathcal{F}$ with normal bundle $N$ is said to have degree $d$ if $N\simeq{\mathcal{O}}_{X}(d+2)$, where ${\mathcal{O}}_{X}(1)$ denotes the hyperplane bundle. In this case, we denote the space of degree-$d$ foliations by ${\sf Fol}(X,d)$. For $X={\mathbb{P}^{n}}$, $n\geq 3$, the cases $d=0,1,2$ are completely classified, and partial results exist for $d=3$ (see [Deserti-Cerveau], [JP-Jouanoulou79], [LNC-96], and [CLP-2022]). For more general smooth projective varieties, few results describe the structure of foliations of low degree. When $d=0$, classifications are available for certain homogeneous varieties [BFM, kuster2025], for hypersurfaces [Fig23], and for smooth complete intersections [ACM-dist].
@@
Let $X$ be a smooth complete intersection in $\mathbb{P}^{n+s}$} of type $(d_{1},\ldots,d_{s})$, such that either
  1. (1)
    $X$ is a hypersurface, i.e. $s=1$, $n\geq 4$, and $c_{1}(TX)\geq 2$; or
  2. (2)
    $d_{j}\geq 3$ for all $j=1,\dots,s$, $n\geq 4$, and $c_{1}(TX)\geq 2$; or
  3. (3)
    $d_{j}\geq 4$ for all $j=1,\dots,s$ and $n\geq 3$.
Then ${\sf Fol}(X,1)$ has exactly two irreducible components.
The general elements in these components are defined by logarithmic forms (see §2.2). In particular, all foliations in these spaces arise as restrictions of foliations on the ambient projective space (see Proposition 3.1). When $X$ is the smooth three-dimensional quadric, it is shown in [LPT13] that ${\sf Fol}(X,1)$ possesses an additional irreducible component, consisting of rigid foliations arising from actions of the affine Lie algebra on the quadric, totaling three components.
Among hypersurfaces, the smooth 3-dimensional quadric is the only exceptional case. Theorem A applies to all smooth hypersurfaces either of degree $\deg(X)\leq n$ and dimension $n\geq 4$; or degree $\deg(X)\geq 4$ and dimension $n\geq 3$. We complement this result by showing that it also holds for cubic threefolds:
@@
To prove Theorem A, we focus on varieties covered by lines. We study the deformations of lines along the leaves of a foliation, as introduced in [LPT-def]. First, we show that a general free line is not invariant under any foliation (see Proposition 4.7).
@@
Let $X\subset{\mathbb{P}^{N}}$ be a smooth projective variety of dimension $n\geq 3$ with $\operatorname{Pic}(X)=\mathbb{Z}$, covered by lines. If $\mathcal{F}$ is a codimension-one holomorphic foliation on $X$ of degree $d$ such that either $d=1$ or
@@
Specializing to degree-one foliations on smooth complete intersections with $c_{1}(TX)\geq 2$, we conclude that either the foliation is defined by a closed rational 1-form without codimension-one zeros, or it admits a codimension-two subfoliation of degree zero (see Corollary 4.9). In the latter case, we prove that the foliation is the pullback of a degree-one foliation on the projective plane, and is therefore again defined by a closed rational 1-form without codimension-one zeros. The last item follows from Theorem 3.2. This completes the proof of Theorem A.
@@ 2. Holomorphic foliations
Let $X$ be a smooth projective complex manifold. A distribution $\mathcal{F}$ on $X$ is the data of a saturated subsheaf ${T_{\mathcal{F}}}\subset{TX}$, often called the tangent sheaf of $\mathcal{F}$. The quotient ${N_{\mathcal{F}}}=TX/{T_{\mathcal{F}}}$ is a torsion-free sheaf called the normal sheaf. The singular locus $\operatorname{sing}(\mathcal{F})$ is the locus where $N_{\mathcal{F}}$} is not locally free; since $N_{\mathcal{F}}$} is torsion free, $\operatorname{codim}\operatorname{sing}(\mathcal{F})\geq 2$. Over the smooth locus $X\setminus\operatorname{sing}(\mathcal{F})$, these sheaves fit in a short exact sequence:
$0\longrightarrow{T_{\mathcal{F}}}\longrightarrow{TX}\longrightarrow{N_{\mathcal{F}}}\longrightarrow 0.$
@@
The map ${TX}\to{N_{\mathcal{F}}}$ is defined by contraction with a twisted $p$-form $\omega_{\mathcal{F}}\in H^{0}(X,\Omega^{p}_{X}\otimes\det({N_{\mathcal{F}}}))$, where $p=\operatorname{rk}{N_{\mathcal{F}}}$. Conversely, given a line bundle $L\in\operatorname{Pic}(X)$, $\omega\in H^{0}(X,\Omega^{p}_{X}\otimes L)$ defines a distribution if and only if it is locally decomposable off the singular set, LDS for short, see [Med]. This condition means that for every point $x\in X\setminus\operatorname{sing}(\omega)$ there exists an open neighborhood $U\ni x$ and $1$-forms $\alpha_{1},\dots,\alpha_{p}$ such that
@@
The main objective of our work is to describe the irreducible components of this space for $p=1$, $X\subset{\mathbb{P}^{N}}$ a smooth complete intersection, and specified $L$. Any line bundle on $\mathbb{P}^{n}$} is isomorphic to ${\mathcal{O}}_{\mathbb{P}^{n}}(k)$ for some $k\in{\mathbb{Z}}$. A distribution has degree $d$ if it is defined by $\omega\in H^{0}(\Omega_{\mathbb{P}^{n}}^{p}(d+p+1))$. This follows Lins Neto’s convention of the degree of a foliation. In general, if $X$ is a smooth projective variety such that $\operatorname{Pic}(X)={\mathbb{Z}}$, we define the degree of a foliation $\mathcal{F}$ determined by a section in $H^{0}(X,\Omega_{X}^{p}(d+p+1))$ as $\deg(\mathcal{F}):=d$. In this case, for simplicity, we will denote ${\sf Fol}(X,d):={\sf Fol}^{1}(X,\mathcal{O}_{X}(d+2))$. We are particularly interested in smooth complete intersections.
@@
In [ACM-dist, TeseMateus, Fig23] the authors study distributions and foliations on smooth complete intersections satisfying an extra hypothesis: $X=V(f_{1},\dots,f_{k})$ with $\deg(f_{1})\leq\dots\leq\deg(f_{k})$ such that $X_{i}=V(f_{1},\dots,f_{i})$ is smooth for every $i$. There are smooth complete intersections without this property; for instance, $X=V(f_{1},f_{2})$ where $\deg(f_{1})<\deg(f_{2})$ and $V(f_{1})$ is singular. However, one can verify that the arguments in these papers depend uniquely on the existence of a chain of smooth complete intersection $X_{1}\supset\cdots\supset X_{k}=X$, regardless of the degrees. Therefore, in light of Lemma 2.1, their results apply to smooth complete intersections in general.
@@
Note further that a smooth complete intersection only admits distributions and foliations of nonnegative degree. Indeed, $H^{0}(X,\Omega_{X}^{p}(d+p+1))=0$ for $d<0$ and $X$ a smooth complete intersection, see [ACM-dist, Lemma 5.17].
@@ 2.1. Algebraically integrable foliations
A foliation on $X$ is algebraically integrable if the leaf through a general point is an algebraic subvariety. In this case, there exists a dominant rational map $\pi\colon X\dashrightarrow Y$ onto a projective variety $Y$ such that the closure of a leaf is an irreducible component of a fiber of $\pi$, see [AD-fano, §3].
@@
Let $X$ be a complex projective manifold of dimension greater than $2$ and with $\operatorname{Pic}(X)\cong{\mathbb{Z}}$. Let $\mathcal{F}$ be a codimension-one foliation on $X$. Suppose that $\mathcal{F}$ is algebraically integrable. Then either:
@@
This follows directly from the arguments in [LPT13, Lemma 3.1 and Theorem 3.2]. ∎
@@
Let $X\subset{\mathbb{P}^{N}}$ be a smooth complete intersection of dimension $n\geq 3$ and $\mathcal{F}$ be a codimension-one holomorphic foliation on $X$ of degree $d$. If $\mathcal{F}$ is algebraically integrable, then
@@
By Proposition 2.2, either $\mathcal{F}$ admits a rational first integral $\rho\colon X\dashrightarrow{\mathbb{P}^{1}}$ with only irreducible fibers or admits a subfoliation $\mathcal{G}$. Suppose we are in the first case. By Lefschetz Hyperplane Theorem, $X$ is simply connected. Then the argument follows the proof of [LPT13, Proposition 3.5]. By [LPT13, Theorem 3.3], $\rho$ admits at most $2$ multiple fibers, which we may assume to be over 0 and $\infty$, and write $\rho^{-1}(0)=pH_{0}$ and $\rho^{-1}(\infty)=qH_{\infty}$. Then $\rho^{*}(\frac{dx}{x}-\frac{dy}{y})$ is a logarithmic $1$-form defining $\mathcal{F}$ with polar divisor $H_{0}+H_{\infty}$ and empty zero divisor. ∎
Notice that the proof of Corollary 2.3 holds for any simply connected manifold $X$ with Picard number $1$. For our purposes, we will only need $X$ to be a smooth complete intersection, especially hypersurfaces.
@@ 2.2. Irreducible components of the space of foliations
where $L_{1}$ and $L_{2}$ are degree one homogeneous polynomials, and thus they are all conjugate to each other (see [Deserti-Cerveau, Proposition 3.1]). Therefore, ${\sf Fol}({\mathbb{P}^{n}},0)$ is isomorphic to the Grassmannian of lines in $\mathbb{P}^{n}$}.
For codimension-one foliations of degree one, Jouanoulou in [JP-Jouanoulou79] proved that ${\sf Fol}({\mathbb{P}^{n}},1)$ has two irreducible components. Both components parameterize logarithmic foliations, which we now describe in more generality.
Let $X$ be any smooth projective complex variety such that $\dim(X)\geq 3$, $\operatorname{Pic}(X)={\mathbb{Z}}$ and $H^{1}(X,{\mathbb{C}})=0$. Given $d_{1},d_{2},\dots,d_{k}$ positive integers and $\lambda_{1},\lambda_{2},\dots,\lambda_{k}$ complex numbers, not all zero, such that $\sum\lambda_{i}\cdot d_{i}=0$, we can take sections $f_{i}\in H^{0}(X,{\mathcal{O}}_{X}(d_{i}))$ to form global sections
@@
Then $\omega$ is integrable and, if its singular locus has codimension greater than one, it determines a foliation known as a logarithmic foliation. Varying $\lambda_{i}$ and $f_{i}$, $i=1,\dots,k$, the forms $\omega$ sweep an irreducible algebraic variety subvariety of ${\sf Fol}(X,d_{1}+\dots+d_{k}-2)$ whose closure is denoted by $\textsf{Log}(X;d_{1},d_{2},\dots,d_{k})$. In [Calvo94], it was established that $\textsf{Log}(X;d_{1},d_{2},\dots,d_{k})$ is an irreducible component of ${\sf Fol}(X,\sum d_{i}-2)$. It is called the logarithmic component. When $k=2$, the foliation defined by this 1-form $\omega=\lambda_{1}f_{2}df_{1}+\lambda_{2}f_{1}df_{2}$ is algebraically integrable; this component is also called a rational component. These varieties can also be defined by the closure of the rational map
@@
where $\mathbb{L}_{2}$ is the set of linear projections of $\mathbb{P}^{n}$} to $\mathbb{P}^{2}$}. Then, in [CL-NE], it was demonstrated that $\textsf{LPB}(d)$ is an irreducible component of ${\sf Fol}({\mathbb{P}^{n}},d)$, and it is called the linear pullback component. It is an elementary exercise to show that $\textsf{LPB}(1)=\textsf{Log}({\mathbb{P}^{n}};1,1,1)$.
In degree two, Cerveau and Lins Neto proved in [LNC-96] that ${\sf Fol}({\mathbb{P}^{n}},2)$ has six irreducible components: a linear pullback component $\textsf{LPB}(2)$; the rational components $\textsf{Log}({\mathbb{P}^{n}};3,1)$ and $\textsf{Log}({\mathbb{P}^{n}};2,2)$; the logarithmic components $\textsf{Log}({\mathbb{P}^{n}};2,1,1)$ and $\textsf{Log}({\mathbb{P}^{n}};1,1,1,1)$; and another irredutible component called exeptional component, which is determined by the closure of the map
@@
$\beta=\frac{1}{x_{0}}(2fdg-3gdf)\in H^{0}({\mathbb{P}^{3}},\Omega^{1}_{{\mathbb{P}^{3}}}(4)),$
for $f=2x_{0}x_{2}-x_{1}^{2}$ and $g=3x_{0}^{2}x_{3}-3x_{0}x_{1}x_{2}+x_{1}^{3}$.
In [CLP-2022], the irreducible components of ${\sf Fol}({\mathbb{P}^{n}},3)$ were partially classified; there are at least 24 irreducible components. Besides the projective space, there is some literature on irreducible components of the space of foliations on, for instance, Fano three-folds [LPT13], homogeneous spaces [BFM, kuster2025], and hypersurfaces [ACM-dist, Fig23].
In the case of smooth projective hypersurfaces $X$ in $\mathbb{P}^{n+1}$}, $n>3$, if
@@
then the restriction of foliations from the ambient $\mathbb{P}^{n+1}$} defines an isomorphism between ${\sf Fol}(X,d)$ and ${\sf Fol}({\mathbb{P}^{n+1}},d)$, see [Fig23, Theorem 3.6]. Therefore, the previous classifications for foliations on projective spaces are valid for these varieties. The problem is then to study low-degree hypersurfaces; for degree-one foliations, we need to analyze hypersurfaces of degree at most three.
In [LPT13], the authors proved that for $X\subset{\mathbb{P}^{4}}$ a smooth quadric hypersurface, ${\sf Fol}(X,1)$ has three irreducible components; only two are restrictions from foliations on $\mathbb{P}^{4}$}. In this work, we prove that, for hypersurfaces, this is the only case where ${\sf Fol}(X,1)$ has more than two irreducible components.